# FindingDory: A Benchmark to Evaluate Memory in Embodied Agents

### v1

@@ Abstract
Large vision-language models have recently demonstrated impressive performance in planning and control tasks, driving interest in their application to real-world robotics. However, deploying these models for reasoning in embodied contexts is limited by their ability to incorporate long-term experience collected across multiple days and represented by vast collections of images. Current VLMs typically struggle to process more than a few hundred images concurrently, highlighting the need for more efficient mechanisms to handle long-term memory in embodied settings. To effectively evaluate these models for long-horizon control, a benchmark must specifically target scenarios where memory is crucial for success. Existing long-video QA benchmarks overlook embodied challenges like object manipulation and navigation, which demand low-level skills and fine-grained reasoning over past interactions. Moreover, effective memory integration in embodied agents involves both recalling relevant historical information and executing actions based on that information, making it essential to study these aspects together rather than in isolation. In this work, we introduce a new benchmark for long-range embodied tasks in the Habitat simulator. This benchmark evaluates memory-based capabilities across 60 tasks requiring sustained engagement and contextual awareness in an environment. The tasks can also be procedurally extended to longer and more challenging versions, enabling scalable evaluation of memory and reasoning. We also present baselines that integrate state-of-the-art VLMs with low level navigation policies, assessing their performance on these memory-intensive tasks and highlight areas for improvement. Our dataset and code will be made available here.
Figure 1: FindingDory: We introduce a benchmark designed to evaluate an agent’s memory capabilities by constructing scenarios that necessitate reasoning over previously collected experiences, including those involving various interactions. In these scenarios, an agent is provided with logs from prior interactions in indoor environments and tasked with completing navigation, pick, and place instructions. The figure illustrates a sample episode: first, a privileged agent, with full access to the environment, executes a set of randomly sampled navigation and manipulation tasks. Subsequently, the memory-evaluating agent is introduced into the same environment and tasked with completing instructions that require leveraging the logged experiences of the privileged agent to determine where to navigate or which object to retrieve.
**footnotetext: Equal Contribution
@@ 1 Introduction
Memory is a fundamental capability for intelligent agents that allows them to recall past experiences, adapt to dynamic environments, and make informed decisions over extended timescales. In humans and animals, it plays a crucial role in navigation, reasoning, and goal-directed behavior. As we strive to develop more capable embodied systems, equipping them with robust memory mechanisms is essential — particularly in settings where agents must process high-dimensional multimodal inputs to interact with their surroundings. The ability to store and retrieve relevant information is the key to unlocking sophisticated behaviors, from household assistance to autonomous exploration.
@@
Recent advances in Vision-Language Models (VLMs) have significantly improved high-level reasoning and planning for embodied tasks. These models leverage large-scale multimodal training to exhibit strong scene understanding and action-guided perception. However, most existing VLM applications focus on short-term or static tasks, such as image captioning or Visual Question Answering (VQA), which do not require sustained memory usage. Extending these models to long-horizon embodied control introduces new challenges: unlike VQA, which often involves processing a single image or short video clip, embodied agents must integrate observations over time, recall past experiences, and act accordingly within their environment.
@@
While recent efforts have sought to extend long-context understanding in VLMs—primarily through video-based QA or long-document comprehension [36, 44, 21]—these approaches fail to fully capture the complexities of long-term memory in decision-making. Many existing QA benchmarks either rely on multiple-choice formats, which enable guessing, or require extensive human annotation to curate question-answer pairs that require models to attend to long video sequences. In contrast, we propose embodied memory tasks—such as “navigating to a soft toy that you did not rearrange yesterday” or “find the second object picked up earlier”—that demand fine-grained recall and multi-hop reasoning over past observations. These tasks provide a more rigorous test of memory capabilities, requiring agents to track and respond to environmental changes rather than passively perceive static information.
@@
In this paper, we introduce FindingDory, a benchmark designed to evaluate long-horizon memory in highly photorealistic simulated environments. Our benchmark consists of 60 diverse navigation tasks, which we categorize along various axes of memory requirements, that require both long-range temporal and spatial reasoning while discouraging heuristic shortcuts or random guessing. To ensure that performance is directly tied to memory utilization rather than simple perception, we curate scenarios where agents must recall past interactions to succeed. Additionally, our benchmark features dynamic environments - where agent interactions modify the scene - making it essential to reason over changing contexts. Furthermore, FindingDory is procedurally extensible, allowing tasks to scale in complexity as vision-language models improve, ensuring continued relevance as embodied AI progresses.
@@
Unlike standard QA datasets that rely on subjective human annotations—such as manually selected keyframes or hand-crafted question-answer pairs—our benchmark leverages photorealistic simulation to enable automatic, objective evaluation of memory-based navigation. These tasks go beyond static recall since they require agents to perform complex spatio-temporal reasoning over past interactions. For instance, tasks such as “navigate to any object that you interacted with yesterday” demand not only accurate memory retrieval, but also strategic decision-making to identify and navigate to the nearest valid target from where the agent is currently positioned. In FindingDory, we introduce metrics that quantify task completion efficiency and not just successful recall. Our key contributions are:
@@ 2 Related Work
MemoryMaze [27] One-hot $1$ 500-1000 steps
MemoryGym [28] One-hot $3$ 128-512 steps (Extensible)
MultiON [42] One-hot $1$ 500 steps
SIF [25] Language $3$ $240$ (val/test) ✗ (Uses semantic map)
OpenEQA (Active) [23] Language $1600$ 75 steps
GoatBench [16] Language + Image $1800-3600$ (val) 500 steps (Static Env)
Excalibur [52] Language $11$ 21 Requires exploration
@@
2.1 Video QA benchmarks
Recent analyses of video QA benchmarks [24, 7] show that many tasks can be solved by attending to key frames, not full video sequences, suggesting that current benchmarks may favor keyframe selection over evaluating long-term memory. Long-context evaluations for foundation models also focus on “needle-in-a-haystack” tasks, evaluating retrieval of specific details rather than integrating information across many segments [50]. While recent video benchmarks feature more challenging questions [24, 19, 41, 11, 36, 35, 18, 30, 47, 51, 23], they adopt multiple-choice QA formats which are not suitable for evaluating fine-grained decision-making over long histories. They may also allow models to succeed through guessing and are not extensible since they require substantial human annotation effort.
@@ 2.2 Embodied AI benchmarks
Various benchmarks have been proposed to evaluate multi-modal language models in embodied settings where agents must leverage their experience in environments to improve future decision-making [42]. Many focus on embodied question answering (EQA), such as HM-EQA [31], which requires active exploration, and Open-EQA [23], which tests VLMs in a passive setting with pre-collected episode histories. Notably, Open-EQA found that a blind LLM baseline—one that ignores visual input—still performed well, suggesting that multiple-choice evaluation can introduce biases and reduce the need for genuine memory recall. Other benchmarks study the situated instruction following setting where agents must gather information to interpret ambiguous instructions [32, 20, 25]. However, many allow agents to sidestep memory challenges by precomputing relevant objects and locations based on recent history, effectively turning potentially memory-intensive tasks into structured instruction-following problems. Memory-Maze and Memory-Gym [27, 28] focus on partially observable decision-making with history-based state modeling, but their simplistic visuals and small state spaces limit their relevance to complex real-world memory settings.
@@
    Scope of semantic reasoning needed over memory: Distinguishes between benchmarks that require agents to recall diverse semantic information versus a narrow set of attributes such as object locations.
@@
    Memory horizon length: Adapting the definition from [26], this measures the temporal scope over which memory must be maintained for optimal task performance. Unlike static benchmarks requiring localized frame recall, dynamic environments demand reasoning over events distributed across extended observation sequences. This metric becomes difficult to quantify when tasks require substantial exploration phases.
@@ 3 FindingDory
In this section, we describe the design and construction of our embodied memory benchmark, FindingDory. We first outline the benchmark structure, then detail the constituent tasks and simulation setup.
3.1 Benchmark Overview
@@
The goal of FindingDory is to evaluate an agent’s ability to retain and utilize past experiences to efficiently complete future tasks. Unlike standard navigation or perception-based evaluations, our benchmark explicitly tests memory by structuring tasks around past interactions rather than only immediate observations. To achieve this, we design a two-phase setup consisting of an experience collection phase, where scripted oracle agents autonomously gather experiences in the environment, followed by an interaction phase, where they are evaluated on their ability to recall and act upon that history. This decomposition allows for a controlled and standardized evaluation of memory, ensuring that the agent’s ability to remember past interactions is isolated from its ability to explore effectively.
The experience collection phase introduces dynamism into the environment by having the agent interact with and rearrange different objects. This ensures that the history contains meaningful state changes, allowing us to construct tasks that require spatio-temporal reasoning. Using oracle information from the simulator, the agent picks up objects from one location and places them onto designated receptacles (e.g., a table, chair, or bed). Episodes vary in length, spanning 400 to 3500 frames, with 2 to 11 object pick-and-place interactions (further details in Sec. B.2).
Once the experience collection phase ends, the interaction phase begins. Here, a task is sampled from a predefined set of templated instructions, requiring the agent to recall and act upon the state of the environment from the previous phase. The agent is provided with access to its full history, including recorded images, pose, and past actions, and is tested on its ability to integrate long-term memory into decision-making.
Due to the photorealistic visuals and object diversity of the Habitat simulator, FindingDory provides a practical and transferable setting for evaluating large-scale vision-language models. Since many VLMs are pretrained on real-world images, they can readily interpret frames from this environment—enabling realistic memory evaluations in simulation, where long-horizon, multimodal interaction data can be collected at scale. This makes FindingDory especially suitable for studying how current foundation models integrate memory into decision-making.
@@
Table 2: Task Categories with Example Prompts and their Challenges. See Appendix F for detailed examples.
{tblr}
width = colspec = X[0.04,l,m]X[0.9,l,m]X[3.0,l,m]X[2.1,l,m], hline1,2,13, row1 = font=, bg=lightgray!30, roweven = bg=lightgray!10, cell12,3,4 = font=, c, column1 = colsep=6pt, column2 = colsep=1pt, column3 = colsep=1pt, column4 = colsep=1pt, & Task Category Example Instructions Memory Requirement
[…]
@@
[…]
@@
[…]
@@
3.2 Tasks
Our benchmark is designed to evaluate a broad spectrum of memory challenges faced by embodied agents, moving beyond narrow tasks that could be solved using simple heuristics or non-generalizable strategies. We define several templated tasks that evaluate an agent’s ability to integrate and recall different forms of memory across spatial, temporal, and semantic dimensions. Some tasks require recalling specific objects or receptacles, tracking when (e.g., temporal ordering) and how (e.g., interaction type: picked vs. placed). Others demand awareness of room-level context (e.g., rooms visited, farthest location) or semantic attributes (e.g., shape, color, material, functionality). We also introduce time-usage memory tasks, such as identifying the room that took the longest to explore or the object with the longest rearrangement duration. Task instructions are generated by randomly sampling scene-specific entities, and the full list of 60 templates is shown in Table LABEL:tab:all_tasks.
A smaller subset of attribute-based evaluations is also included to assess whether agents can flexibly reason over visual and semantic memory. These tasks involve prompts that refer to objects indirectly via properties observed during prior interaction (e.g., “go to the striped object you interacted with previously”). This setup helps differentiate methods capable of recalling fine-grained visual details from those relying solely on static representations, such as semantic maps [8] or textual summaries [13], which often struggle to flexibly reconstruct such information.
Single vs. Multiple Goals. The majority of tasks in FindingDory involve single-goal instructions, where the agent must reason over its experience to identify a target and navigate to it. A smaller subset contains multi-goal tasks that require reaching multiple destinations, sometimes in a specific order. Single-goal tasks can be equally challenging, often requiring agents to recall what was not interacted with or to trace specific object movements through rearrangement states—demanding more than simple recall from static frames.
Extensibility through Procedural Generation. Our task suite is procedurally extensible, allowing difficulty to scale with changes in the experience collection phase. Increasing the number of pick-and-place interactions, for instance, expands the temporal reasoning space and creates more challenging memory dependencies.
Task Categorization. We group tasks into interpretable categories to analyze performance trends. Table 2 presents these categories along with representative instructions and the memory challenges they embody. To our knowledge, FindingDory offers the most visually complex and extensible suite of embodied memory tasks to date, covering a wide range of interaction-driven and long-horizon reasoning scenarios.
@@ 3.3 Implementation details
Simulation Setup and Dataset Composition. We build FindingDory using the Habitat simulator [33, 37] and scenes from the Habitat Synthetic Scenes Dataset (HSSD)[15]. Our benchmark uses 107 scenes from the training split and 30 from the validation split, resulting in 1,478 training and 100 validation episodes. We populate these scenes with the same object sets used in OVMM[48], combining assets from AI2Thor [17], Amazon-Berkeley Objects [9], Google Scanned Objects [10], and HSSD [15].
Across these episodes, we include 839/247 object instances spanning 84/72 distinct categories in the train/val splits, respectively. Objects are rearranged during the experience collection phase to induce temporally grounded memory challenges. Additionally, agents interact with 17/16 distinct receptacle categories in the train/val splits (e.g., tables, chairs, beds). To support attribute-based tasks (see Table 2), we generate semantic descriptions of object instances using GPT-4o. We prompt the model with multi-view images of each object to extract attributes such as shape, color, material, and functionality. These descriptions are used to populate the object attribute task templates (5 total), which are limited to the validation set and undergo manual quality review.
In total, our benchmark includes 82,174 training and 5,876 validation task instances. Details on the procedural generation pipeline for episodes and instructions are provided in Appendix B.
Agent Setup. We adopt the Hello Robot Stretch embodiment [14] in our simulation. The agent is equipped with RGB and depth sensors (resolution $640\times 480$) and a GPS+Compass sensor. The agent can execute one of four discrete navigation actions: MOVE_FORWARD ($0.25$m), TURN_LEFT / TURN_RIGHT (by $10^{\circ}$) and a STOP action. During the experience collection phase, the agent can also manipulate its arm joints (extension, lift, yaw, pitch, roll) and activate a manipulation_mode for pick-and-place actions, which rotates the base ($90^{\circ}$) and directs the head camera toward the gripper.
PDDL-based Task Verification. Since we operate with a large and diverse set of instruction templates (see Table 2), we leverage the PDDL [2] specification system as instantiated in [38]. For each instruction template, we create a template goal condition that uses various predicate expressions such as is_robot_at_object(x, "OBJECT" or is_robot_at_receptacle(x, "RECEPTACLE" where x is some placeholder variable. These placeholder variables are sampled based on various constraints such as target_color or interaction_order that are defined through the template. Together, the placeholder constraints and goal conditions enable us to create scalable, automated verification routines for each task in FindingDory.

### v2

@@ Abstract
Vision-Language models (VLMs) have recently demonstrated impressive performance in planning and control tasks, driving interest in their application to robotics. Yet their deployment in embodied settings remains limited by the challenge of incorporating long-term experience, often spanning multiple days and represented by vast image collections. Current VLMs typically handle only a few hundred images at once, underscoring the need for more efficient mechanisms to manage long-term memory in embodied contexts. To meaningfully evaluate these models for long-horizon control, a benchmark must target scenarios where memory is essential. Existing long-video QA benchmarks neglect embodied challenges like object manipulation and navigation, which require low-level skills and fine-grained reasoning over past interactions. Moreover, effective memory integration in embodied agents involves both recalling relevant historical information and executing actions based on that information, making it essential to study these aspects together. In this work, we introduce FindingDory, a new benchmark for long-range embodied tasks in the Habitat simulator. FindingDory evaluates memory-centric capabilities across 60 tasks requiring sustained engagement and contextual awareness in an environment. The tasks can also be procedurally extended to longer and more challenging versions, enabling scalable evaluation of memory and reasoning. We further present baselines that integrate state-of-the-art closed-source and fine-tuned open-source VLMs with low-level navigation policies, assessing their performance on these memory-intensive tasks and highlighting key areas for improvement. [^1] Website: https://findingdory-benchmark.github.io/
@@ 1 Introduction
Figure 1: FindingDory: We propose a benchmark for evaluating agents’ memory by constructing scenarios that require reasoning over previously collected experiences. In these scenarios, an agent receives logs of prior interactions in indoor environments and must complete navigation, pick, and place instructions. The figure shows a sample episode: first, a privileged agent with full environment access executes randomly sampled navigation and manipulation tasks. Then, the memory-evaluating agent is placed in the same environment and tasked with completing instructions that rely on the privileged agent’s logs to decide where to navigate or which object to retrieve.
@@
Memory is a fundamental capability for intelligent agents that allows them to recall past experiences, adapt to dynamic environments, and make informed decisions over extended timescales. In humans and animals, it plays a crucial role in navigation, reasoning, and goal-directed behavior. As we strive to develop more capable embodied systems, equipping them with robust memory mechanisms is essential, particularly in settings where agents must process high-dimensional multimodal inputs to interact with their surroundings. The ability to store and retrieve relevant information is the key to unlocking sophisticated behaviors, from household assistance to autonomous exploration.
@@
Recent advances in Vision-Language Models (VLMs) have significantly improved high-level reasoning and planning for embodied tasks. These models leverage large-scale multimodal training to exhibit strong scene understanding and action-guided perception. However, most existing VLM applications focus on short-term or static tasks, such as image captioning or Visual Question Answering (VQA), which do not require sustained memory usage. Extending these models to long-horizon embodied control introduces new challenges: embodied agents must integrate observations over time, recall past experiences, and act accordingly within their environment.
@@
While recent efforts have sought to extend long-context understanding in VLMs, primarily through video-based QA or long-document comprehension (Song et al., 2023; Xiong et al., 2025; Ma et al., 2024), these approaches fail to fully capture the complexities of long-term memory in decision-making. Many existing QA benchmarks either rely on multiple-choice formats, which enable guessing, or require extensive human annotation to curate question-answer pairs that require models to attend to long video sequences. In contrast, we propose embodied memory tasks—such as “navigating to a soft toy that you did not rearrange yesterday” or “find the second object picked up earlier” that demand fine-grained recall and multi-hop reasoning over past observations. These tasks provide a more rigorous test of memory capabilities, requiring agents to track and respond to environmental changes rather than passively perceive static information.
@@
In this paper, we introduce FindingDory, a benchmark designed to evaluate long-horizon memory in highly photorealistic simulated environments. Our benchmark consists of 60 diverse navigation tasks, which we categorize along various axes of memory requirements, that require both long-range temporal and spatial reasoning while discouraging heuristic shortcuts or random guessing. To ensure that performance is directly tied to memory utilization rather than simple perception, we curate scenarios where agents must recall past interactions to succeed. Additionally, our benchmark features dynamic environments - where agents modify the scene - making it essential to reason over changing contexts. Furthermore, FindingDory is procedurally extensible, allowing tasks to scale in complexity as VLMs improve, ensuring continued relevance as embodied AI progresses.
@@
Unlike standard QA datasets that rely on subjective human annotations, our benchmark leverages photorealistic simulation to enable automatic evaluation of memory-based navigation. These tasks go beyond static recall since they require agents to perform complex spatio-temporal reasoning over past interactions. For instance, tasks such as “navigate to any object that you interacted with yesterday” demand not only accurate memory retrieval, but also strategic decision-making to identify and navigate to the nearest valid target from where the agent is currently positioned. In FindingDory, we introduce metrics that quantify task completion efficiency and not just successful recall.
Our key contributions are:
@@ 2 Related Work
2.1 Video QA benchmarks
Recent analyses of video QA benchmarks (Mangalam et al., 2023; Buch et al., 2022) show that many tasks can be solved by attending to key frames, not full video sequences, suggesting that current benchmarks may favor keyframe selection over evaluating long-term memory.
@@ 2.1 Video QA benchmarks
MemoryMaze (Pašukonis et al., 2023) One-hot 11 500-1000 steps
MemoryGym (Pleines et al., 2023) One-hot 33 128-512 steps (Extensible)
MultiON (Wani et al., 2020) One-hot 11 500 steps
SIF (Min et al., 2024) Language 33 $240$ (val/test) ✗ (Uses semantic map)
OpenEQA (Active) (Majumdar et al., 2024) Language $1600$ 75 steps
GoatBench (Khanna* et al., 2024) Language + Image 180036001800-3600 (val) 500 steps (Static Env)
Excalibur (Zhu et al., 2023) Language $11$ 21 Requires exploration
@@
Long-context evaluations for foundation models also focus on “needle-in-a-haystack” tasks, evaluating retrieval of specific details rather than integrating information across many segments (Zhang et al., 2024). While recent video benchmarks feature more challenging questions (Mangalam et al., 2023; Li et al., 2023; Wang et al., 2024; Fu et al., 2024; Song et al., 2023; 2024; Li et al., 2023; Rawal et al., 2024; Yang et al., 2024; Zhou et al., 2024; Majumdar et al., 2024), they adopt multiple-choice QA formats which are not suitable for evaluating fine-grained decision-making over long histories. They may also allow models to succeed through guessing and are not extensible since they require substantial human annotation effort.
@@ 2.2 Embodied AI benchmarks
Various benchmarks have been proposed to evaluate multi-modal language models in embodied settings where agents must leverage their experience in environments to improve future decision-making (Wani et al., 2020). Many focus on embodied question answering (EQA), such as HM-EQA (Ren et al., 2024), which requires active exploration, and Open-EQA (Majumdar et al., 2024), which tests VLMs in a passive setting with pre-collected episode histories.
Notably, Open-EQA found that a blind LLM baseline—one that ignores visual input—still performed well, suggesting that multiple-choice evaluation can introduce biases and reduce the need for genuine memory recall. Other benchmarks study the situated instruction following setting where agents must gather information to interpret ambiguous instructions (Sashank Dorbala et al., 2024; Ma et al., 2023; Min et al., 2024). However, many allow agents to sidestep memory challenges by precomputing relevant objects and locations based on recent history, effectively turning potentially memory-intensive tasks into structured instruction-following problems. Memory-Maze and Memory-Gym (Pašukonis et al., 2023; Pleines et al., 2023) focus on partially observable decision-making with history-based state modeling, but their simplistic visuals and small state spaces limit their relevance to complex real-world memory settings.
@@
    Scope of reasoning needed over memory: Distinguishes between benchmarks that require recalling diverse semantic information v.s. a narrow set of attributes (e.g. object locations).
@@
    Memory horizon length: Adapting the definition from Ni et al. (2023), this measures the temporal scope over which memory must be maintained for optimal task performance. Unlike static benchmarks requiring localized frame recall, dynamic environments demand reasoning over events distributed across extended observation sequences. This metric becomes difficult to quantify when tasks require substantial exploration phases.
@@ 3 FindingDory
In this section, we introduce our embodied memory benchmark, FindingDory, built on top of the Habitat simulator Savva et al. (2019). Habitat offers photorealistic visuals, diverse household scenes, and object-rich environments, making it a practical and transferable testbed for evaluating memory-augmented embodied agents. Since many vision-language models (VLMs) are pretrained on real-world image distributions, they can directly interpret observations from Habitat, enabling realistic and scalable evaluation of long-term decision-making.
We first describe the overall structure of FindingDory and the suite of tasks it comprises, then provide implementation details.
3.1 Benchmark Design and Tasks
@@
The goal of FindingDory is to evaluate an agent’s ability to retain and utilize past experiences to efficiently complete future tasks. To achieve this, we design a two-phase setup: an experience collection phase, where scripted oracle agents autonomously gather experiences in the environment, and an interaction phase, where agents are evaluated on their ability to recall and act upon that history. This decomposition isolates memory from exploration, enabling standardized evaluation.
During experience collection, oracle agents pick up and place objects onto designated receptacles (e.g., a table, chair, or bed), introducing meaningful state changes into the environment. Episodes span 400–3500 frames with 2–11 pick-and-place interactions (see section B.2). Once collection ends, the interaction phase begins: the agent is given access to its recorded history (images, pose, past actions) and tasked with following instructions that require integrating long-term memory into decision-making.
Tasks are instantiated from a set of templated instructions designed to probe diverse memory challenges. These cover spatial, temporal, and semantic reasoning: e.g., recalling which object was moved, when it was moved, or how it was interacted with; identifying rooms visited or farthest locations; and reasoning over attributes such as shape, color, or material. We also introduce time-usage memory tasks (e.g., longest room exploration, object with longest manipulation). A smaller subset of attribute-based tasks involves referring to objects indirectly via observed properties (e.g., “go to the striped object you interacted with previously”), helping distinguish methods that can flexibly reconstruct fine-grained visual details from those relying on static summaries. In total, 60 templates are used to generate scene-specific task instances (see LABEL:tab:all_tasks).
Single vs. Multiple Goals. Most tasks involve single-goal instructions, but some require reaching multiple destinations in sequence. Even single-goal tasks can be challenging, demanding agents recall what was not interacted with or trace object movements across rearrangement states.
Extensibility. The task suite is procedurally extensible: adjusting the complexity of experience collection (e.g., number of interactions) scales the temporal reasoning space and memory dependencies. We further group tasks into interpretable categories for analysis (see table 2), making FindingDory one of the most visually complex and extensible embodied memory benchmarks to date.
@@
Table 2: Task categories with example prompts and challenges. See appendix G for more examples.
Task Category Example Instructions Memory Requirement
Spatial
Object Recall Navigate to a {category}. Object was previously seen or needs exploration.
@@
Temporal
Interaction Order Navigate to the object you interacted with immediately after {object_category}. Sequence of objects interactions.
@@
Multi-Goal
Unordered Revisitation Revisit all the receptacles you picked objects from yesterday. Past locations of multiple objects and best visitation order.
@@ 3.2 Implementation details
Simulation Setup and Dataset Composition. We build FindingDory using the Habitat simulator (Savva et al., 2019; Szot et al., 2021) and scenes from the Habitat Synthetic Scenes Dataset (HSSD) (Khanna* et al., 2023). Our benchmark uses 107 scenes from the training split and 30 from the validation split, with 1,478 training and 100 validation episodes. We populate these scenes with the same object sets used in OVMM (Yenamandra et al., 2023). Across these episodes, we include 839/247 object instances spanning 84/72 distinct categories in the train/val splits, respectively. Objects are rearranged during the experience collection phase to induce temporally grounded memory challenges. Additionally, agents interact with 17/16 distinct receptacle categories in the train/val splits (e.g., tables, chairs, beds). To support attribute-based tasks (see table 2), we generate semantic descriptions of object instances using GPT-4o. We prompt the model with multi-view images of each object to extract attributes such as shape, color, material, and functionality. These descriptions are used to populate the object attribute task templates (5 total), which are limited to the validation set and undergo manual quality review. In total, our benchmark includes 79,213 training and 5,876 validation task instances. More details in appendix B.
Agent Setup. We adopt the Hello Robot Stretch embodiment (Kemp et al., 2022) in our simulation. The agent is equipped with RGB and depth sensors (resolution $640\times 480$) and a GPS+Compass sensor. The agent can execute one of four discrete navigation actions: MOVE_FORWARD (0.250.25m), TURN_LEFT / TURN_RIGHT (by $10^{\circ}$) and a STOP action. During the experience collection phase, the agent can also manipulate its arm joints (extension, lift, yaw, pitch, roll) and activate a manipulation_mode for pick-and-place actions, which rotates the base ($90^{\circ}$) and directs the head camera toward the gripper.
Procedural Generation and Verification using PDDL. We operate over a large set of diverse task instructions (see  table 2) by encoding them into templates using the PDDL specification system  (Aeronautiques et al., 1998; Szot et al., 2024). Specifically, each template specifies entities (objects and receptacles) with properties such as interaction order, target color, and start/goal receptacle. These entities are combined using predicates that map to pythonic functions which can be evaluated against any simulator state. Goal conditions compose these predicates with quantifiers (e.g., EXISTS, FOR_ALL_UNORDERED) to express complex logical relationships between multiple predicates. The defined instructions templates are procedurally expanded by sampling valid entity bindings based on the entity properties which serve as constraints. The bindings are finally substituted into the goal conditions to generate the FindingDory dataset and enable scalable, automated verification for each episode without hand-coding low-level rules.